OpenStudy (anonymous):

collect like terms 16a+4b-14a-6b=?

OpenStudy (anonymous):

2a-2b

OpenStudy (anonymous):

2a -2b

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

add the numbers with a and then add the numbers with b

OpenStudy (anonymous):

16a-14a = 2a 4b -6b = -2b

OpenStudy (anonymous):

16a - 14a = 2a 4b - 6b = -2b so 2a-2b